# A Unified Framework for Iris Anti-Spoofing: Introducing IrisGeneral Dataset and Masked-MoE Method

### v1

@@ Abstract
Iris recognition is widely used in high-security scenarios due to its stability and distinctiveness. However, the acquisition of iris images typically requires near-infrared illumination and near-infrared band filters, leading to significant and consistent differences in imaging across devices. This underscores the importance of developing cross-domain capabilities in iris anti-spoofing methods. Despite this need, there is no dataset available that comprehensively evaluates the generalization ability of the iris anti-spoofing task. To address this gap, we propose the IrisGeneral dataset, which includes 10 subsets, belonging to 7 databases, published by 4 institutions, collected with 6 types of devices. IrisGeneral is designed with three protocols, aimed at evaluating average performance, cross-racial generalization, and cross-device generalization of iris anti-spoofing models. To tackle the challenge of integrating multiple sub-datasets in IrisGeneral, we employ multiple parameter sets to learn from the various subsets. Specifically, we utilize the Mixture of Experts (MoE) to fit complex data distributions using multiple sub-neural networks. To further enhance the generalization capabilities, we introduce a novel method Masked-MoE (MMoE). It randomly masks a portion of tokens for some experts and requires their outputs to be similar to the unmasked experts, which improves the generalization ability and effectively mitigates the overfitting issue produced by MoE. We selected ResNet50, VIT-B/16, CLIP, and FLIP as representative models and benchmarked them on the IrisGeneral dataset. Experimental results demonstrate that our proposed MMoE with CLIP achieves the best performance on IrisGeneral.
Introduction
The iris is inside the eye and has large differences between each human, which is stable and distinguishable (Agarwal and Jalal 2021), making it widely used in high-security scenarios. Thus it is necessary to ensure the security of the iris recognition system because it always relates to the property and critical areas (Nigam, Vatsa, and Singh 2015; Perala 2017; Gent 2023). Iris recognition systems are similar to other kinds of biometric detection systems that are weak to spoofing such as paper prints, screens, contact lenses, and plastic eyes (Yadav, Chen, and Ross 2019, 2020).
Figure 1: The unified framework for iris anti-spoofing is shown in this Figure. Including the IrisGeneral dataset and the Masked-MoE method.
DataBase Dataset Devices. Nickname
@@ Introduction
Clarkson 2015 for LivDet-Iris 2015 Clarkson2015Dalsa DALSA Db 700 / 873 378 / 558 2509
@@
IFVE IFVEAI1000 AI1000 F 20000 / 20000 5000 / 5000 50000
Notre Dame 2013 for LivDet-Iris 2013 ND2013LG4000 LG 4000 Ga 2000 / 1000 800 / 400 4200
@@
Notre Dame 2017 for LivDet-Iris 2017 ND2017LG4000 LG 4000 Gb 600 / 600 900 / 900 3000
@@
Table 1: This Table shows the scale of the sub-datasets employed in the IrisGeneral dataset, including the numbers of real and fake samples in both training and testing sets. The first letters of the nicknames denote the collection devices while the second letters denote the order to distinguish.
Among those, contact lenses are complicated to detect because they are worn on real eyeballs and can deceive liveness detection systems (Bowyer and Doyle 2014; Galbally et al. 2012). They are transparent with some textures, so the iris texture and the contact lens texture are intertwined, making them difficult to classify.
@@
Some previous work has been performed to combat contact lens attacks, such as 2D Gabor wavelet-based iris recognition (Daugman 2009), and 4DCycle-GAN (Zou et al. 2018). They trained on specific datasets and showed good results, but may not be able to cope with the problem once the acquisition equipment is changed. The near-infrared fill light bands, camera filters, collection distances, and ambient light, make a lot of impact on iris acquisition which requires near-infrared presentation (Xiao et al. 2013). The image-imaging differences between each iris acquisition device are obvious and fixed. This means that the cross-domain ability of the iris anti-spoofing methods is necessary, thus the security could be maintained more effectively. However, there are no suitable iris datasets to train and evaluate iris anti-spoofing models’ generalization on cross-domain adaptation (Yadav and Ross 2024). To address that, we proposed this IrisGeneral dataset, which includes 10 subsets, belonging to 7 databases, published by 4 institutions, and collected with 6 types of devices. It is designed with three protocols, aimed at evaluating average performance, cross-racial generalization, and cross-device generalization of iris anti-spoofing models. Section IrisGeneral Dataset provides the details of this dataset.
@@
The IrisGeneral brings the challenge of integrating multiple sub-datasets with different racial and devices. To tackle it, we employ multiple parameter sets to learn from the various subsets by using the Mixture of Experts (MoE) (Jacobs et al. 1991) to fit complex data distributions with multiple sub-neural networks. However, some previous works have found that sparse MoE would bring overfitting problems into models (Shazeer et al. 2017; Chen et al. 2023; Elbayad, Sun, and Bhosale 2022; Lepikhin et al. 2020). At the same time, overfitting is also the main challenge of the iris anti-spoofing task, which is caused by its high structural homogeneity, leading to the differences between live iris and fake iris mainly in their fine textures. Thus, alleviating the overfitting and improving the generalization of the model is necessary. To address that, we introduce a novel Masked-MoE (MMoE) method to enhance the generalization capabilities. It has two contributions: Mask Mechanism to alleviate overfitting, and Cosine Distance Loss to further improve the model’s generalization. Specifically, the Mask Mechanism masks parts of the input tokens for several experts and allows one expert to see all tokens with complete information. Then the Cosine Distance Loss requests the output of masked experts to be similar to the unmasked experts as shown in Figure 1. It is inspired by MAE (He et al. 2022), and we refer to this masking method to improve the generalization to deal with the inherent overfitting problem of MoE. Section Proposed Method provides the details of our MMoE method. To evaluate our contributions, we selected ResNet50, VIT-B/16, CLIP, and FLIP as representative models and benchmarked them on the IrisGeneral dataset. Experimental results show that our proposed MMoE with CLIP has strong generalization abilities and achieves the best performance on both three protocols, section Experiment provides more details.
The main contributions in this paper are shown as follows:
  • We proposed a unified framework for Iris Anti-Spoofing, including IrisGeneral dataset and Masked-MoE method.
@@
  • Dataset: The IrisGeneral dataset is proposed with three protocols that could respectively evaluate the average, cross-racial, and cross-device performance of the iris-antispoofing task.
@@
  • Method: The Masked-MoE (MMoE) method is proposed with the Mask Mechanism and the Cosine Distance Loss, which efficiently alleviates the problem of MoE method overfitting and improves the models’ generalization.
Related Work
Iris Anti-Spoofing
Detecting contact lenses is a complicated task of iris recognition system. Many previous research focuses on this, including traditional and deep-learning-based methods (Yadav and Ross 2024; Kumar, Lamba, and Jangra 2020). Wei et al. (Wei et al. 2008) proposed a method to detect the sharpness of the edge of the iris. He et al. (He, An, and Shi 2007) conducted classification through a feature vector and support vector machine based on GLCM. J. Komulainen et al. (Komulainen, Hadid, and Pietikäinen 2014) studied the generalized cases of software-based device-independent detection of contact lens. Many studies deal the iris anti-spoofing by neural networks such as CNN (LeCun et al. 1998) or generative network-based methods (Goodfellow et al. 2014; Kingma and Welling 2013), and etc. For example,  (Yadav et al. 2014a), and (Tinsley, Flynn, and Czajka 2024; Boyd et al. 2021).
Cross-domain Adaption
Cross-domain adaption measures the models’ ability to transfer the knowledge learned in the source domain to the target domain. Two main directions to improve are as follows. One is introducing an adversarial discriminator (Tzeng et al. 2017) to bring the source domain and the target domain closer. The other direction is extracting the features of source and target domain with one neural network (Ren et al. 2020). A loss function is designed to measure the distance between the two domains. There also some previous works which proposed effective methods to improve the adaption, for example, DAN (Long et al. 2015), MMD (Yadav et al. 2014b), and JAN (Long et al. 2017).
Figure 2: Some image samples of the IrisGeneral dataset. Include 10 subsets with Real and Fake iris images.
IrisGeneral Dataset
Subsets Detail
There are 10 subsets, belonging to 7 databases, published by 4 institutions, collected with 6 types of devices in the IrisGeneral Dataset. Table 1 lists the information of each subset. Figure 2 provides some real and fake iris image samples of each subset.
CASIAH100 (Sun et al. 2014) with 6744 samples in total, collected by H100, belonging to database CASIAH100, contains 5398 samples in its training set and 1346 samples in its testing set. We call this dataset H in short.
Clarkson2013Dalsa (Yambay et al. 2017) with 1356 samples in total, collected by DALSA, belonging to database Clarkson 2013 for LivDet-Iris 2013, contains 670 samples in its training set and 686 samples in its testing set. We call this dataset Da in short.
Clarkson2015Dalsa (Yambay et al. 2017) with 2509 samples in total, collected by DALSA, belonging to database Clarkson 2015 for LivDet-Iris 2015, contains 1573 samples in its training set and 936 samples in its testing set. We call this dataset Db in short.
Clarkson2015LG2200 (Yambay et al. 2017) with 1944 samples in total, collected by LG2200, belonging to database Clarkson 2015 for LivDet-Iris 2015, contains 990 samples in its training set and 954 samples in its testing set. We call this dataset La in short.
Clarkson2017LG2200 (Yambay et al. 2017) with 5841 samples in total, collected by LG2200, belonging to the database Clarkson 2017 dataset for LivDet-Iris 2017, contains 3591 samples in its training set and 2250 samples in its testing set. We call this dataset Lb in short.
IFVEAI1000 (Zhang et al. 2021) with 50000 samples in total, collected by AI1000, belonging to database IFVE, contains 40000 samples in its training set and 10000 samples in its testing set. And we call this dataset F in short.
ND2013LG4000 (Doyle and Bowyer 2014)with 4200 samples in total, collected by LG4000, belonging to database Notre Dame 2013 for LivDet-Iris 2013, contains 3000 samples in its training set and 1200 samples in its testing set. We call this dataset Ga in short.
ND2013AD100 (Doyle and Bowyer 2014)with 900 samples in total, collected by AD100sensor, belonging to database Notre Dame 2013 for LivDet-Iris 2013, contains 600 samples in its training set and 300 samples in its testing set. We call this dataset Aa in short.
ND2017LG4000 (Yambay et al. 2017) with 3000 samples in total, collected by LG4000, belonging to database Notre Dame 2017 for LivDet-Iris 2017, contains 1200 samples in its training set and 1800 samples in its testing set. We call this dataset Gb in short.
ND2017-CDAD100 (Yambay et al. 2017) with 1800 samples in total, collected by AD100sensor, belonging to database Notre Dame 2017 for LivDet-Iris 2017, contains 0 samples in its training set and 1800 samples in its testing set. We call this dataset Ab in short.
Dataset
P1
P2
P3.1
P3.2
P3.3
P3.4
P3.5
P3.6
[…]

### v2

# A Unified Framework for Iris Anti-Spoofing: Introducing Iris Anti-Spoofing Cross-Domain-Testing Protocol and Masked-MoE Method

@@ Abstract
Iris recognition is widely used in high-security scenarios due to its stability and distinctiveness. However, iris images captured by different devices exhibit certain and device-related consistent differences, which has a greater impact on the classification algorithm for anti-spoofing. The iris of various races would also affect the classification, causing the risk of identity theft. So it is necessary to improve the cross-domain capabilities of the iris anti-spoofing (IAS) methods to enable it more robust in facing different races and devices. However, there is no existing protocol that is comprehensively available. To address this gap, we propose an Iris Anti-Spoofing Cross-Domain-Testing (IAS-CDT) Protocol, which involves 10 datasets, belonging to 7 databases, published by 4 institutions, and collected with 6 different devices. It contains three sub-protocols hierarchically, aimed at evaluating average performance, cross-racial generalization, and cross-device generalization of IAS models. Moreover, to address the cross-device generalization challenge brought by the IAS-CDT Protocol, we employ multiple model parameter sets to learn from the multiple sub-datasets. Specifically, we utilize the Mixture of Experts (MoE) to fit complex data distributions using multiple sub-neural networks. To further enhance the generalization capabilities, we propose a novel method Masked-MoE (MMoE), which randomly masks a portion of tokens for some experts and requires their outputs to be similar to the unmasked experts, which can effectively mitigate the overfitting issue of MoE. For the evaluation, we selected ResNet50, VIT-B/16, CLIP, and FLIP as representative models and benchmarked them under the proposed IAS-CDT Protocol. The results show our proposed method MMoE with CLIP achieves the best performance, which demonstrates its generalization performance and its training effectiveness on our proposed IAS-CDT Protocol.
Index Terms:
Iris Anti-Spoofing, Mixture of Experts, CLIP, Cross-Domain-Testing Protocol.
I Introduction
Figure 1: This figure shows the unified framework for iris anti-spoofing. Including the Iris Anti-Spoofing Cross-Domain-Testing (IAS-CDT) Protocol and Masked-MoE (MMoE) Method.
The iris has large differences between each human, which is stable and distinguishable [1, 2], making it widely used in high-security scenarios. Thus, it is necessary to ensure the security of the iris recognition system because it always relates to the property and critical areas [3, 4, 5]. Iris recognition systems are similar to other kinds of biometric detection systems that are vulnerable to spoofing, such as paper prints, screens, contact lenses, and plastic eyes [6, 7]. Among those, contact lenses are complicated to detect because they are worn on real eyeballs and can fool the liveness detection algorithems [8, 9], which harm the security of the iris recognition systems [10, 11, 12, 13]. There are various types of contact lenses, most of which are transparent with some textures in public datasets. It is intertwined with the original textures of the iris, making the identity difficult to classify [14].
DataBase Dataset Devices Nickname
@@ I Introduction
Clarkson 2015 for LivDet-Iris 2015  [16] Clarkson2015Dalsa DALSA Db 700 / 873 378 / 558 2509
@@
IFVE [17] IFVEAI1000 AI1000 F 20000 / 20000 5000 / 5000 50000
Notre Dame 2013 for LivDet-Iris 2013  [18] ND2013LG4000 LG4000 Ga 2000 / 1000 800 / 400 4200
@@
Notre Dame 2017 for LivDet-Iris 2017  [16] ND2017LG4000 LG4000 Gb 600 / 600 900 / 900 3000
@@
TABLE I: This Table shows the scale of the sub-datasets employed in the Iris Anti-Spoofing Cross-Domain-Testing (IAS-CDT) Protocol, including the numbers of real and fake samples in both training and testing sets. The first letters of the nicknames denote the collection devices, while the second letters denote the order to distinguish.
Extensive previous works combat contact lens attacks, primarily categorized into hardware-based, software-based, and challenge-response approaches [14]. Both hardware-based and challenge-response approaches need extra devices when detecting. The hardware-based one needs special sensors, in addition to standard iris sensors, to measure the special characteristics of eyes, in contrast, the challenge-response one utilizes extra devices to stimulate the eyes and detect the spoofing samples by comparing the captured eyes’ responses. Unlike these two, software-based approaches merely use static features extracted from samples acquired by standard iris sensors. For traditional software-based methods, the study mainly focused on features, they designed different extractors to get the distinctive features from iris images and used them for classification. There are several represent works such as the sharpness of edges [13], gray level value [19], Local Binary Pattern (LBP) [20, 21, 22], Binary Statistical Image Features (BSIF) [23], Scale-Invariant Feature Transform (SIFT) [24, 15], a fusion of 2D and 3D features [25], Local Phase Quantization (LPQ) [26], etc. Deep-learning-based methods mainly amend the structure of the models, compared to previous methods, instead of modifying the extraction and design of features, which simplifies the feature extraction process and allows models to learn the critical differences between real samples and the spoofing ones by updating the parameters themselves, the advantage of deep-learning-based methods. For example, the Attention Learning [27, 28] methods aim to fuse multi-scale features, which could improve the performance of the CNN framework, make the models concentrate more on the global features, and alleviate the overfitting caused by CNN focusing on local features. The Adversarial Learning [6, 29, 30] aims to generate unseen iris samples for anti-spoofing models to improve their generalization, and the convolutional structure [31] aims to extract the critical features using neural networks.
Nowadays, the development of Large Language Models (LLMs) provides a novel way to deal with these Biometric Anti-Spoofing problems, which could use the pre-trained models to extract the features of the samples. Some novel works show good performance on the Face Anti-Spoofing (FAS) task with LLMs, for example, the FLIP [32] proposed by Srivatsan et al. combines the visual transformers and language guidance based on the CLIP [33], and fine-tunes it with different strategies to form three variants: FLIP-Vision (FLIP-V), FLIP-Image-Text Similarity (FLIP-IT), and FLIP-Multimodal-Contrastive-Learning (FLIP-MCL). The Class-Free Prompt Learning (CFPL) [34] proposed by Liu et al. introduces two lightweight transformers, Content Q-Former (CQF) and Style Q-Former (SQF), to learn different semantic prompts based on content and style features. It leverages the large-scale vision-language model CLIP to dynamically adjust the classifier’s weights, significantly enhancing the generalization of extracted features. Zhang et al. introduced a multimodal fine-grained CLIP (MFCLIP) [35], which designed an innovative Sample Pair Attention (SPA) module, which can adaptively highlight relevant negative sample pairs, optimizing cross-modal feature alignment, greatly improving the generalization ability for detecting diffusion-generated facial fakes.
@@
Though the deep-learning-based methods have good results in experiments, they are trained on specific datasets, which makes it hard to cope with the situation once the collection devices are changed. The previous study proves that the image-imaging differences between each iris collection device are obvious and fixed. The near-infrared fill light bands, camera filters, collection distances, and ambient light, make a lot of impact on iris acquisition which requires near-infrared presentation [36], which could bring negative impacts on iris anti-spoofing (IAS) models in practice. This means that the cross-domain ability of the IAS methods is necessary, thus the security could be maintained more effectively. However, there are no suitable datasets or protocols to train and evaluate IAS models’ generalization on cross-domain adaptation [37].
To address this gap, we propose an Iris Anti-Spoofing Cross-Domain-Testing (IAS-CDT) Protocol, which involves 10 datasets, belonging to 7 databases, published by 4 institutions, and collected with 6 different devices. It contains three sub-protocols hierarchically. Protocol-1 evaluates the model’s average performance. Protocol-2 evaluates the model’s generalization performance of cross-racials. Protocol-3 evaluates the model’s generalization performance of cross-devices. Section III provides more details of this IAS-CDT protocol.
The proposal of the IAS-CDT Protocol brings the challenge of integrating multiple sub-datasets with different races and devices. To tackle it, we employ multiple parameter sets to learn from the various subsets by using the Mixture of Experts (MoEs) [38] to fit complex data distributions with multiple sub-neural networks. The concept of MoEs has evolved over at least 30 years [39], originating with the work of Jacobs et al. in 1991 [38]. Initially, MoEs were introduced as a supervised learning framework known as ”adaptive mixtures of local experts,” where multiple independent networks collaboratively processed subsets of training samples. More recently, the application of MoEs has gained prominence through the work of Shazeer et al., who introduced the Sparsely-Gated Mixture-of-Experts Layer [40]. This layer enabled sparse gating and token-level processing, paving the way for scalability in large language models (LLMs). Based on this, Lepikhin, Dmitry, et al. proposed GShard [41] represents a significant milestone as the first to implement MoE within a Transformer framework, enhancing efficiency in large-scale training, improving LLMs model performance on sparse tasks. Since then, more and more MoE-related works have been proposed, such as V-MoE [42], ST-MoE [43], Switch Transformer [44], LIMoE [45], DeepseekMoE [46], MH-MoE [47] and etc..
@@
However, some previous works have found that sparse MoE would bring overfitting problems into models [40, 48, 49, 41]. It may be caused by the low utilization of parameters in sparse MoEs, which means that single or several experts integrate multi-tasks and other experts learn redundant knowledge because only parts of the experts deal with the input, and the parameters of the unused experts are not updated. At the same time, overfitting is also the main challenge of the iris anti-spoofing task, which is caused by its high structural homogeneity, leading to the differences between live iris and fake iris mainly in their fine textures. Thus, alleviating the overfitting and improving the generalization of the model is necessary. In that case, we introduce a novel Masked-MoE (MMoE) method to enhance the generalization capabilities and reduce the passive overfitting impacts brought by MoE. It has two contributions: Mask Mechanism to alleviate overfitting, and Cosine Distance Loss to further improve the model’s generalization. Specifically, the Mask Mechanism masks parts of the input tokens for several experts and allows one expert to see all tokens with complete information. Then the Cosine Distance Loss requests the output of masked experts to be similar to the unmasked experts as shown in Figure 1. It is inspired by MAE [50], and we refer to this masking method to improve the generalization to deal with the inherent overfitting problem of MoE. The Mask Mechanism could reduce the overfitting problem by dropping out a reasonable rate of the input iris features, while the Cosine Distance improves the model’s generalization by driving different Experts in MoE to adapt unknown knowledge by closing the cosine distance between masked features and unmasked features. Section Proposed Method provides the details of our MMoE method.
As we mentioned before, the development of LLMs provides a novel way to deal with Biometric Anti-Spoofing problems. Since CLIP [33] has already performed its powerful and competitive ability in the FAS field, for example UniAttackDetection [51], La-SoftMoE [52], CPL-CLIP [53], we choose it as the backbone of our proposed Iris Anti-Spoofing model. CLIP [33] is the visual-linguistic dual-modality pre-trained model. It has two encoders, an image encoder and a text encoder, the contrastive learning loss is employed to drive the training of the two encoders mapping the feature of different modalities to the same latent space with similar information to close the distance of image embedding and text embedding.
We selected ResNet50 [54], ViT-B/16 [55], CLIP [33], and FLIP [32] as representative models to evaluate our contributions and benchmarked them on the IrisGeneral dataset. Experimental results show that our proposed MMoE with CLIP has strong generalization abilities and achieves the best performance on both three protocols, section V provides more details.
To sum up, the main contributions in this paper are shown as follows:
  • We proposed a unified framework for Iris Anti-Spoofing, including an Iris Anti-Spoofing Cross-Domain-Testing (IAS-CDT) Protocol and a Masked-MoE Method. To address the gap of no suitable datasets or protocols to train and evaluate IAS models’ generalization on cross-domain adaptation. And improve the competition of iris recognition systems in the anti-spoofing area.
@@
  • Protocol: Iris Anti-Spoofing Cross-Domain-Testing. The IAS-CDT Protocol is proposed with three sub-protocols that could respectively evaluate the average, cross-racial, and cross-device performance of the iris-antispoofing task.
@@
  • Method: Masked-MoE.The MMoE method is proposed with the Mask Mechanism and the Cosine Distance Loss, which efficiently alleviates the problem of MoE method overfitting and improves the models’ generalization.
II Related Work
This section introduces several previous works on Iris Anti-Spoofing (IAS), Cross-domain Adaptation, Cross-domain Adaptation in Biometric Anti-Spoofing, and Mixture-of-Experts (MoEs).
II-A Iris Anti-Spoofing (IAS)
The evolution of iris anti-spoofing (IAS) methods has experienced three distinct phases, each marked by different feature engineering paradigms. As shown in Fig 2, the phases are Handcrafted Feature Engineering, Local Descriptor Fusion, and Deep-Learning-Based methods respectively.
Figure 2: This figure shows the development of the iris anti-spoofing (IAS) methods, mainly separated into three phases.
The first phase, Handcrafted Feature Engineering (2008-2014), saw some IAS research focus on designing manual texture descriptors for SVM-based classification, while other research explored new classifiers to replace SVM due to its costly computation. Within SVM-based research, there are still many different types of approaches that stood out. There are some GLCM-based methods. For instance, Wei et al. [13] proposed a texture analysis approach for fake iris detection, incorporating features like iris edge sharpness, Iris-Texton features, and GLCM-based features. Alonso et al. [56] presented an iris anti-spoofing method using GLCM for feature extraction paired with an SVM classifier, which saved time and reduced complexity. Methods based on the Wavelet Transform also performed well on IAS field. He et al. [10] introduced an anti-spoofing method based on Wavelet Packet Transform (WPT) and SVM, effectively capturing distinctive features between fake and genuine irises. Zhang et al. [20] developed an iris anti-spoofing method using SIFT descriptors and weighted LBP (w-LBP), with SVM as the classifier. Another dominant approach was the Hierarchical-Visual-Codebook-based method. Sun et al. [15] proposed an iris image classification method based on the Hierarchical Visual Codebook (HVC),which included iris image preprocessing, low-level visual feature extraction, statistical iris image representation based on the HVC model, and iris image classification.
Some research in this phase also focused on classifier innovation beyond SVM. He et al. [22] proposed an anti-spoofing decision model based on LBP and Adaboost learning, which effectively differentiated genuine and fake iris images, showing strong performance against specific spoofing methods like contact lenses.
The second phase, Local Descriptor Fusion (2014-2015), saw researchers refine feature granularity for spoofing iris detection and explore models’ generalization ability on cross-device tasks. Gupta et al. [21] introduced an iris anti-spoofing approach using SVM as the classifier with traditional feature extraction methods like LBP, Histogram of Oriented Gradients (HOG), and GIST. Raghavendra et al. [23] developed a Presentation Attack Detection (PAD) method for face and iris anti-spoofing, employing linear SVM for decision-making.
The third phase, Deep-Learning-Based methods (2015-present), divides IAS research into hybrid architectures and fully deep-learning-based methods. Hybrid architectures blended deep learning with manual features. For example, Chen et al. [28] combined the advantages of the attention mechanism and CNN, proposing a novel method to detect the detection of iris presentation attacks, named Attention-Guided PAD (AG-PAD). Fang et al. [27] proposed an IAS method, called Attention-based Pixel-level Binary Supervision (A-PBS), based on attention mechanism and CNN as well. Wu et al. [57] proposed an unsupervised domain adaptation transfer learning model to enhance the precision and generalization of iris anti-spoofing, leveraging contrastive learning to enhance characteristic characteristics of features.
While these methods have performed well in IAS, there’s a lack of large-scale or wide-ranging generalization evaluation. The development of multi-modal pre-trained models and LLMs offers new ways to improve IAS model performance and generalization. Thus, we propose an IAS model based on CLIP and MoE (a key LLM component), evaluating its performance and generalization on our proposed IAS-CDT Protocol, which includes many datasets suitable for generalization evaluation.
II-B Cross-domain Adaption
The evolution of domain adaptation methods in pattern recognition has been marked by two major paradigm shifts: Adversarial Learning Dominance and Feature Space Unification.
The Adversarial Learning Dominance paradigm was initiated by adversarial training methods, which align domain distributions through minimax optimization. For example, Tzeng et al. [58] proposed the Adversarial Discriminative Domain Adaptation (ADDA), an unsupervised domain generalization method that combines adversarial learning with discriminative feature learning. A domain discriminator is trained to minimize the distance between the source and target domains, enabling direct application of the source classification model to target representations. This combination of discriminative modeling and adversarial loss addresses limitations of traditional methods in handling large-scale domain shifts. Long et al. [59] introduced Joint Adaptation Networks (JAN) to address data distribution disparities between source and target domains. Unlike methods focusing only on marginal distribution alignment, JAN emphasizes joint distribution adaptation across multiple layers. It has demonstrated excellent classification accuracy on various benchmark datasets, validating its effectiveness in deep transfer learning.
Feature Space Unification approaches, unlike adversarial learning methods that enhance model domain generalization, with the rise of large models bypass adversarial training by constructing domain-invariant feature spaces. For example, Long et al. [60] introduced a Deep Adaptation Network (DAN) to improve the transferability of deep neural networks in domain adaptation. Ren et al. [61] proposed Target Domain-Specific Classifier Learning (TSCDA) to address knowledge transfer challenges when source and target domain label spaces are inconsistent. TSCDA includes two interdependent modules: Selective Weighted MMD (SWMMD) for partial feature alignment and PEers Assisted Learning (PEAL). It minimizes the prediction discrepancy between the target and auxiliary classifiers, promoting collaborative learning among multiple target-specific classifiers and enhancing their discriminative power, effectively mitigating negative transfer and classifier bias in partial domain adaptation.
Although Adversarial Learning Dominance methods are theoretically flexible and universal, their training challenges make feature space unification methods more practical for real-world applications.
II-C Cross-domain Adaption in Biometric Anti-Spoofing
Domain generalization presents a significant challenge for biometric anti-spoofing systems, as performance often degrades due to cross-domain discrepancies, including sensor variations, shifts in attack types, modality differences, and other factors.
Early approaches utilized deep learning architectures, particularly CNNs, to establish baseline generalization capabilities. Silva et al. [25] employing a CNN to construct deep image representations and a fully connected layer with softmax regression for classification, and demonstrated substantial performance improvements in cross-domain tests. Safaa et al. [31] focused on PAD for face and iris modalities, featuring a deep CNN with multi-layer convolutional and pooling architecture, and produced binary classification results through fully connected layers. Boyd et al. [62] presented CYBORG, which integrates human-annotated saliency maps into the loss function. It achieved higher accuracy on unseen samples compared to traditional methods.
To overcome CNN limitations in capturing global dependencies, attention mechanisms emerged as a key advancement. Fang et al. [27] developed the Attention-based Pixel-level Binary Supervision (A-PBS) network, addressing CNN-based PAD method deficiencies like limited local feature capture and overfitting. The A-PBS model outperformed DenseNet and PBS in most cross-database scenarios, validating its generalization capability. Chen et al. [28] proposed an explainable Attention-Guided PAD (AG-PAD) for iris presentation attack detection, integrating attention mechanisms to enhance CNN performance and provide visual interpretability.
The rise of vision-language models spurred a paradigm shift toward multimodal domain adaptation. Zhang et al. [63] introduced CPL-CLIP, that uses dual-branch prompt learning to disentangle cross-modal domain shifts, which enables deployment-robust generalization without target domain knowledge. Fang et al. [64] proposed a visual-language framework (VL-FAS) for cross-domain face anti-spoofing, achieving domain-invariant features through language-driven semantic alignment. Hu et al. [65] presented FGPL, a vision-language framework for cross-domain face anti-spoofing, enabling robust spoofing detection across unseen domains through joint optimization of text-visual cues and domain-invariant features.
Recent innovations emphasize domain-agnostic prompt engineering and frequency-space alignment. Liu et al. [34] proposed Class-Agnostic Prompt Learning (CFPL), advancing cross-domain face anti-spoofing by leveraging vision-language models to disentangle content and style semantics without domain labels. Liu et al. also introduced FM-CLIP [66], which enhances cross-domain adaptation by jointly aligning multimodal features in frequency and text-embedding spaces. Guo et al. [67] developed Style-Conditional Prompt Token Learning (S-CPTL) for cross-domain face anti-spoofing, dynamically generating domain-invariant text prompts conditioned on visual style features to disentangle liveness semantics from domain shifts, advancing generalized biometric defense against domain drift.